OpenStudy (anonymous):

Perimeter of a rectangle is 500 cm. How do you find the maximum area of the rectangle? What is the length and the width?

OpenStudy (ybarrap):

$$ P=2x+2y=5m\\ A=x\times y\\ x>0\\ y>0 $$ We need to maximize A subject to P. $$ 2y = 5-2x\\ y=\cfrac{5}{2}-x\\ A=x\times y=x\times (\cfrac{5}{2}-x)=\cfrac{5}{2}x-x^2\\ $$ Where \(P \) is Perimiter and \(A\) is Area. To maximize A, take derivative and set equal to zero: $$ \cfrac{d A}{dx}=\cfrac{d}{dx}\left(\cfrac{5}{2}x-x^2 \right)=\cfrac{5}{2}-2x=0\\ \implies 2x=\cfrac{5}{2}\implies x=\cfrac{5}{4}\\ \implies A=\cfrac{5}{2}\cfrac{5}{4}-\left(\cfrac{5}{4}\right )^2\\ \qquad \qquad = \cfrac{25}{8}-\cfrac{25}{16}=\cfrac{25\times 2-25}{16}=\cfrac{25}{16} $$ Therefore, the maximum area is \(\cfrac{25}{16}~m^2\).